# Asymptotic Distribution-free Change-point Detection for Modern Data Based on a New Ranking Scheme

### v3

@@ VI-A Kernel and Distance IN Graph CPD
The approach proposed in this paper can also be extended to weights other than ranks in weighting the edges in the similarity graph. For example, we could incorporate kernel value or distance directly to have Kernel IN Graph Change-Point Detection (KING-CPD) and Distance IN Graph Change-Point Detection (DING-CPD) methods. Specifically, we can use the kernel values or (negative) pairwise distances to weight the similarity graph, and the asymptotic property still holds under mild conditions. Let
@@
Besides, Theorem 2 also holds by replacing $R_{ij}$ by $K_{ij}$.
@@ Acknowledgments
The authors were supported in part by the NSF awards DMS-1848579 and DMS-2311399.

### v4

@@ VI-A Kernel and Distance IN Graph CPD
The proposed approach can be extended to use weights other than ranks in weighting the edges in the similarity graph. For example, we could incorporate kernel value or distance directly to have Kernel IN Graph Change-Point Detection (KING-CPD) and Distance IN Graph Change-Point Detection (DING-CPD) methods. Specifically, we can use the kernel values or (negative) pairwise distances to weight the similarity graph, and the asymptotic property still holds under mild conditions. Let
@@
In addition, Theorem 2 also holds by replacing $R_{ij}$ by $K_{ij}$.
@@ Acknowledgments
The authors would like to thank the associate editor and referees for their constructive comments and suggestions.
Doudou Zhou received his Ph.D. in Statistics from the University of California, Davis, in 2022. From 2022 to 2024, he was a Postdoctoral Researcher at Harvard University. He is currently an Assistant Professor in the Department of Statistics and Data Science at the National University of Singapore. His research focuses on developing theoretical and computational methods for high-dimensional data analysis and machine learning.
Hao Chen received her Ph.D. in Statistics from Stanford University in 2013. Since then, she joined the Department of Statistics, University of California, Davis, as a faculty member, where she is currently an Associate Professor (with tenure). Her research interests include developing practical and robust methods that can deal with various data types, including high-dimensional data, image data, and network data, for hypothesis testing, signal detection, classification, and clustering.